OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = –1.7x + 8.5 17x + 10y = 85 –17x + 10y = 85 –17x + 10y = –85 17x + 10y = –85

jimthompson5910 (jim_thompson5910):

First multiply every term by 10 to get y = –1.7x + 8.5 10*y = 10*(–1.7x) + 10*8.5 10y = -17x + 85

jimthompson5910 (jim_thompson5910):

Then add 17x to both sides 10y = -17x + 85 17x + 10y = 85

OpenStudy (anonymous):

Thank you sooo much! That helped (:

jimthompson5910 (jim_thompson5910):

yw